### v4

# Conductivity of concentrated salt solutions

Abstract
The conductivity of concentrated salt solutions has posed a real puzzle for theories of electrolytes. Despite a quantitative understanding of dilute solutions, an analytical theory for concentrated ones remains a challenge almost a century, although a number of parameters and effects incorporated into theories increases with time. Here we show that the conductivity of univalent salt solutions can be perfectly interpreted using a simple model that relies on a modified mean-field description of electrostatic effects and on a classical approach to calculating colloid electrophoresis. We derive a compact equation, which predicts that the ratio of conductivity at a finite concentration to that at an infinite dilution is the same for all salt, if it is plotted against a product of the harmonic mean of ion hydrodynamic radii and the square root of concentration. Our equation fits very well the data for inorganic salts (up to a few mol/l), although at a very high dilution the relaxation correction seems necessary.
preprint: APS/123-QED
I Introduction
The electrolyte solutions are ionic conductors, thanks to cations and anions formed as a result of dissociation. The electrical conductivity $K$ is one of their most important trait, which is widely used in chemical, biological as well as other applications, and the performance of many (e.g. energy storage) devices depends entirely on the ionic transport [1, 2, 3]. The amount of water in Earth’s mantle is inferred from the conductivity data [4]. Besides, historically, the conductivity is the most important source of information on electrolyte properties (e.g. ion pairing) [5].
The physical origin of ionic conductivity is more or less clear. If an electric field $E$ is applied, ions migrate relative to a solvent by generating an electric current of density $J=KE$ (Ohm’s law). The migration speeds of ions are given by
$V_{\pm}=M_{\pm}E,$ (1)
where $M_{\pm}$ are their mobilities. The current densities induced in an univalent electrolyte are then $J_{\pm}=en_{\infty}V_{\pm}$, where $e$ is an elementary positive charge and $n_{\infty}=n_{\pm}$ is the number density (concentration) of an electrolyte solution. Consequently, the current density of a solution reads $J=J_{+}-J_{-}=eEn_{\infty}M,$ where $M=M_{+}-M_{-}$ is the difference in mobilities of cations and anions, and
$K=en_{\infty}M$ (2)
Thus, the calculations of the conductivity at a given concentration are reduced to those of $M$.
The quantitative understanding of ion mobilities is a challenging problem that has been addressed over nearly a century and by many groups, which is often termed a central issue of chemical physics [6, 7]. The simplest expression can be derived by postulating the Stokes resistance to the ion propulsion
$M_{\pm}\simeq M_{\pm}^{0}\simeq\pm\dfrac{e}{6\pi\eta R_{\pm}},$ (3)
where $\eta$ is the dynamic viscosity of the solvent and $R_{\pm}$ are the hydrodynamic radii of the cation and anion. $M_{\pm}^{0}$ given by Eq.(3) and termed the mobilities at an infinite dilution do not depend on salt concentration. In this model
$M\simeq M_{0}\simeq\dfrac{e}{3\pi\eta R_{h}}\,\mathrm{and}\,K\simeq K_{0}\simeq\dfrac{e^{2}n_{\infty}}{3\pi\eta R_{h}},$ (4)
where $\eta$ is the dynamic viscosity of a solvent and $R_{h}=2(R_{+}R_{-})/(R_{+}+R_{-})$ is the harmonic mean of hydrodynamic radii. However, experiments on a conductivity show that $M$ and $K$ are generally smaller than predicted by (4) and this discrepancy augments on increasing salt concentration [8, 1]. Chemists have long used this fact to infer a degree of dissociation (or an ion pairing). Physicists might view this simply as unreliability of hydrodynamic arguments based on the Stokes force.
There is a large literature describing attempts to provide a satisfactory theory of electrolyte conductivity. One of the first systematic treatments of the influence of $n_{\infty}$ on conductivity was contained in a remarkable paper by Onsager [9] who clarified a mechanism for the conductivity reduction, which involved ionic electrophoresis and relaxation, and showed that at a very high dilution the correction to $M_{0}$ depends on the square root of the concentration. In efforts to better understand the connection between the ion mobilities and salt concentration many authors extended the theory, but failed to come to grips with the conductivity at high salt [10]. However, most electrolyte solutions in nature and various applications are concentrated. Thus, salt concentration in human blood plasma is about 0.15 mol/l, in the Atlantic Ocean it is $\simeq 0.6$ mol/l, Li-ion batteries and supercapacitor electrolytes are usually of concentration $1-2$ mol/l, reference electrolytes of pH-meters and glass micropipette electrodes - of concentration 3 mol/l.
Recent attempts at improvements on the theory has been focused on a more sophisticated description of electrostatic and hydrodynamic effects. Various techniques, such as the mean spherical approximation [11, 6], mode-coupling theory [12, 13], density functional approach [12, 14], and more [15, 16, 17, 7, 18] have been employed. These publications involve additional parameters and contributions, and mostly rely on numerical calculations. This makes them difficult to use and limits prediction capabilities, but simple analytical formulas applied for concentrated solutions have not yet been obtained, although there have been some attempts to deduce them from microscopic theories. Useful analytical solution is known [19], but expressed in terms of special functions and can hardly be called simple. Avni et al. [14] derived an approximate conductivity expression, but do not recognize its inapplicability at molar concentrations due to non physical behavior of electrophoretic and relaxation terms [20]. Recently we suggested a completely different line to attack the ionic electrophoresis problem [21]. Assuming cations and anions of equal size, we have carried calculations of an electrophoretic term and derived an equation for a corresponding reduction in relative conductivity, $K/K_{0}$, that applies even at a very low dilution. However, this paper has not given sufficient attention to the relaxation issue, which is traditionally invoked in the conductivity theories. It is also, in practice, unlikely that both cations and anions have exactly the same size, so the status of our theory for real salt solutions remains somewhat obscure.
The present paper extends and generalizes the prior analysis [21] to solutions composed of inorganic ions of unequal radius $R_{+}\neq R_{-}$. We derive a simple equation for an electrophoretic contribution to conductivity valid up to a few mol/l and argue that relative conductivities, $K/K_{0}$, of all inorganic salts plotted against $R_{h}\sqrt{n_{\infty}}$ would collapse into a single curve. This conclusion is supported by providing a comparison with data for several standard salts, which shows that our theory is in excellent agreement with experiment up to concentrations of a few molars. We also demonstrate that although the relaxation effect affects a small decrement to a relative conductivity at high dilution, it becomes negligibly small in concentrated solutions, where the reduction in $K/K_{0}$ is significant.
II General considerations
We consider a bulk 1:1 salt solution assuming that the description of its global static properties can be restricted to a mean-field Poisson-Boltzmann theory. The dimensionless electrostatic potential around each ion $\phi(z)=e\Phi(z)/(k_{B}T)$, where $k_{B}$ is the Boltzmann constant and $T$ is a temperature, represents a continuous function that depends on all other ions. The Debye screening length of a solution, $\lambda_{D}=\left(8\pi\ell_{B}n_{\infty}\right)^{-1/2}$, is defined as usually with the Bjerrum length, $\ell_{B}=\dfrac{e^{2}}{\varepsilon k_{B}T}$, where $\varepsilon$ is the solvent permittivity. By analyzing the experimental data it is more convenient to use the concentration $c_{\infty}[\rm{mol/l}]$, which is related to $n_{\infty}[\rm{m^{-3}}]$ as $n_{\infty}\simeq N_{A}\times 10^{3}\times c_{\infty}$, where $N_{A}$ is Avogadro’s number. The Bjerrum length of water at $T\simeq 298$ K is equal to about $0.7$ nm leading to a useful formula for 1:1 electrolyte
$\lambda_{D}[\rm{nm}]\simeq\frac{0.305[\rm{nm}]}{\sqrt{c_{\infty}[\rm{mol/l}]}}$ (5)
Thus upon increasing $c_{\infty}$ from $10^{-5}$ to $3$ mol/l the screening length is reduced from about 100 down to 0.18 nm.
Figure 1: Sketch of the cation and anion with hydrodynamic radii $R_{+}$ and $R_{-}$ in the bulk electrolyte solution characterized by $\lambda_{D}$. The cation propels with the speed $V_{+}$ in the direction of the electric field $E$. Its zeta potential $\zeta_{+}$ (or dimensionless mobility) is positive. The anion of negative $\zeta_{-}$ migrates against the field with the velocity $V_{-}$.
Table 1: Some typical values of hydrodynamic radii of ions [22], they harmonic means and ratios.
Salt $R_{+}$, nm $R_{-}$, nm $R_{h}$, nm N
KBr 0.1295 0.1179 0.123 1.1
NaCl 0.184 0.1245 0.148 1.48
LiI 0.238 0.1135 0.154 2.1
We recall that the mean-field theory neglects correlations and finite sizes of ions. By contrast, to describe the dynamic response to an external field, the ion hydrodynamic radii $R_{\pm}$ should be taken into account. Inorganic ions have hydrodynamic radii from 0.1 to 0.3 nm [22], and we present their values for some univalent electrolytes in Table 1, together with harmonic means $R_{h}=2R_{+}R_{-}/(R_{+}+R_{-})$ and $N=R_{+}/R_{-}$. These apparent contradictions indicate that in reality, ions are neither point-like nor particles of finite size, exhibiting the properties of the former or the latter only depending on the circumstances. This can be referred to as the point-particle duality of ions (by analogy to the familiar concept of wave-particle duality of electrons in quantum mechanics). Such a duality can be introduced, if an electrolyte solution is modelled as a colloidal suspension of so-called central ions representing isolating particles of radius $R_{\pm}$ with a constant surface charge density ±$\pm e/(4\pi R_{\pm}^{2})$ [21]. Such a sphere induces the same outer field as a point charge ±$\pm e$ located at its center would, and the salt-dependent surface potential $\phi_{s\pm}$ is established self-consistently. By contrast, the electrostatic diffuse layers (EDLs or clouds) are treated on the classical mean-field level and formed by point-like ionic species, thus their extension is defined by $\lambda_{D}$. Two further comments should be made: (i) Since the model prevents point-like EDL ions from penetrating inside the sphere, the need to invoke an additional electrostatic cut-off [14] is removed. (ii) The approach neglects electrostatic correlations inside the cloud, but provides positional correlations between central ions, although on average they are randomly distributed.
Since the electrostatics of ions involves two length scales, it is convenient to introduce their dimensionless radii $\varrho_{\pm}=R_{\pm}/\lambda_{D}$. Note that in the range of $c_{\infty}$ below 3 mol/l we consider here, the values of ϱ±subscriptitalic-$\varrho_{\pm}$ remain smaller than unity or very close to it. Say, if we set $c_{\infty}=2$ mol/l, then for a largest ion (Li+) in Table 1 we obtain $\varrho_{+}\simeq 1$, but for a smallest (I-) we get $\varrho_{-}\simeq 0.5$.
When an electric field $E$ is applied, an electro-osmotic flow around central ions, referred below simply to as ions, is induced. The electroosmosis takes its origin in the EDL, where a tangential electric field generates a force that sets the fluid in motion. The emergence of this flow in turn provides hydrodynamic stresses that cause the propulsion of ions with a velocity given by Eq.(1) as sketched in Fig. 1. The space between clouds of two ions is electro-neutral, and this region is force-free (no flow) provided they do not overlap. Thus, the problem of ionic electrophoresis can be treated on a single-particle level.
We return to the importance of the relaxation later, by focusing now on electrophoresis solely. This contribution to a conductivity is traditionally calculated assuming that the surrounding cloud remains undisturbed during the migration. The electrophoretic mobilities of these ions can be defined using the known expression for an electro-osmotic mobility [23, 24] taken with the opposite sign
$M_{\pm}=\dfrac{\varepsilon Z_{\pm}}{4\pi\eta}=\dfrac{e}{4\pi\eta\ell_{B}}\zeta_{\pm},$ (6)
where $Z_{\pm}$ are the electrokinetic or zeta potentials of cations and anions, and $\zeta_{\pm}=eZ_{\pm}/k_{B}T$. In such a definition the geometry factor is hidden inside $\zeta$. Note that by introducing the dimensional velocity $v_{\pm}=\dfrac{4\pi\eta\ell_{B}}{eE}V_{\pm}$, one can rewrite (1) as $v_{\pm}=\zeta_{\pm},$ which points clearly that the electrophoretic speed of ions is set by their zeta potentials.
III Results and discussion
Zeta potentials of small inorganic ions are given by [21]
$\zeta_{\pm}\simeq\phi_{s\pm}\mathcal{F_{\pm}}\simeq\pm\dfrac{\ell_{B}}{R_{\pm}(1+\varrho_{\pm})}\mathcal{F_{\pm}},$ (7)
where $\mathcal{F_{\pm}}$ are the special functions derived by Henry [25]
±$\mathcal{F_{\pm}}=1-e^{\varrho_{\pm}}\left[5\operatorname{Ei}_{7}(\varrho_{\pm})-2\operatorname{Ei}_{5}(\varrho_{\pm})\right].$ (8)
[…]
±$\mathcal{F_{\pm}}\simeq\dfrac{2}{3}\left[1+\left(\dfrac{\varrho_{\pm}}{4}\right)^{2}\right]$ (9)
[…]
$\zeta_{\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}(1+\varrho_{\pm})}.$ (10)
Clearly, the upper bound for the zeta potential is attained when $\rho_{\pm}\to 0$ (an infinite dilution)
$\zeta_{\pm}\to\zeta_{0\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}}.$ (11)
Figure 2: Zeta potentials $\zeta_{\pm}$ computed for K+, Li+, Cl-, I- (solid curves from top to bottom). Dashed lines are obtained using Eq. (11). Symbols show calculations from Eq. (10).
Figure 2 includes computed zeta potentials of several ions. For these examples we take from Table 1 cations (K+, Li+) and anions (I-, Cl-) of smallest and largest radius. Calculations are made using the first equality in (7) with $\phi_{s\pm}$ found numerically from the nonlinear Poisson-Boltzmann equation and $\mathcal{F_{\pm}}$ calculated from Eq. (8). The straight lines corresponding to $\zeta_{0\pm}$ given by (11) describe perfectly the distinct plateau regions at very low concentrations. On increasing concentration, however, the absolute values of $\zeta_{\pm}$ decrease. Also included are calculations from (10). It can be seen that their fit to numerical data is extremely good in the whole concentration range.
Refer to captionRefer to captionRefer to caption
Figure 3: $K_{\infty}/K_{\infty}^{0}$ as a function of $c_{\infty}$ for KBr, NaCl and LiI aqueous solution calculated from Eqs. (13), (14) and (15) (solid, dash-dotted and dashed curves). Open and filled circles indicate experimental data from [27] and [28], open and filled squares show data from [29] and [30].
From Eq.(6) it follows that
$M\simeq\dfrac{e}{4\pi\eta\ell_{B}}(\zeta^{+}-\zeta^{-})$ (12)
Substituting (10) and performing standard calculations we derive
$\dfrac{M}{M_{0}}\equiv\dfrac{K}{K_{0}}\simeq\dfrac{1}{N+1}\left(\dfrac{1}{1+\varrho_{+}}+\dfrac{N^{2}}{N+\varrho_{+}}\right),$ (13)
where $M_{0}$ and $K_{0}$ are given by Eqs. (4). Figure 3 shows $K/K_{0}$ calculated from Eq. (13) as a function of $c_{\infty}$. The calculations are compared with experimental data for three standard salts [27]. As seen from Table 1, for these examples $N$ varies from 1.1 for KBr, but can be larger than 2 (e.g. for LiI). An overall conclusion from this linear scale plot is that the theoretical curve for $K/K_{0}$ is well consistent with experiment. At large concentrations Eq.(13) predicts either slightly lower (KBr), but still fits data quite well. When $N=1$, we get $\varrho_{h}=R_{h}/\lambda_{D}=2\varrho_{+}/(1+N)=\varrho_{+}$, and Eq. (13) reduces to
$\dfrac{K}{K_{0}}\simeq\dfrac{1}{1+\varrho_{h}}.$ (14)
Eq. (14) predicts a monotonic decrease in relative (electrophoretic) conductivity with $\varrho$ as it should be, and $K/K_{0}\to 0$ as $\varrho\to\infty$. For $\varrho_{h}\ll 1$ it reduces to the famous Onsager formula for the electrophoretic effect
$\dfrac{K}{K_{0}}\simeq 1-\varrho_{h}.$ (15)
The calculation from Eq.(14) is also included in Fig. 3. It can be seen that deviations from Eq. (13) are extremely small. The difference between (13) and (14) is always below a few %. We thus conclude that compact approximate Eq. (14), which is easy to handle, can safely be employed to interpret data for inorganic salts or for a predictive purpose.
It follows from (5) that $\varrho_{h}\simeq 3.379R_{h}\sqrt{c_{\infty}}$ for a fixed $T=298$ K, where $R_{h}$ should be taken in nm. This implies that the values of $K/K_{0}$ for any univalent salt solution plotted against (dimensionless) $R_{h}\sqrt{c_{\infty}}$ should collapse into a single curve. That this is indeed so is demonstrated in Fig. 4, where we plot in a semi-log scale a relative conductivity calculated from (14) vs. $R_{h}\sqrt{c_{\infty}}$ along with the experimental data for a variety of inorganic salts at $T=298$[27, 29, 30, 31, 28]. Besides salts presented in Table 1 we include KCl, LiCl, and LiClO4 using their $R_{h}$ = 0.127, 0.163, and 0.174 nm, correspondingly, to re-scale the concentration [22]. Note that since the theory is justified provided $R_{h}\sqrt{c_{\infty}}$ is below ca. 0.305, we employ for this plot only data for $c_{\infty}\leq 3$ mol/l. One can see that the approximate theory is in a good agreement with the data for all salts, perhaps with only insignificant discrepancy at low dilution that is probably comparable with the experimental error. The discrepancy is always in the direction of smaller $K/K_{0}$ than predicted by Eq. (14), but the data for KBr and KCl show slightly larger $K/K_{0}$. The reason for qualitative differences only for these two salts is unclear and is open for discussion, but it must be remembered that this was a first-order calculation only, and we did not expect it to be very accurate. The curve calculated from Eq. (15) is also included in Fig. 4. Clearly, the data obtained at high salt are irreconcilable with this linear equation.
Figure 4: $K/K_{0}$ as a function of $R_{h}\sqrt{c_{\infty}}$ plotted in a semi-log scale (solid curve) and experimental data for standards salts taken from Refs. [27, 29, 30, 31, 28] shown by filled symbols. Dashed-dotted and dashed curve corresponds to calculations from Eqs. (14) and (15).
Figure 5: A decrement to a relative conductivity, $1-K/K_{0}$, plotted against $R_{h}\sqrt{c_{\infty}}$ in a log-log scale (solid curve). Symbols show the same data as in Fig. 4. Dashed-dotted, dashed and dotted lines are calculated from Eqs. (14),  (15) and (16).
To examine the possible discrepancy from data more closely, the results for salts from Fig. 4 are reproduced in Fig. 5, where a decrement to a relative conductivity, $1-K/K_{0}$, is plotted against $R_{h}\sqrt{c_{\infty}}$ in a log-log scale. It can be seen that in dilute solutions, where the decrease in relative conductivity is so small that was even not discernible on the scale of Fig. 4, the data are slightly above the theory indicating some additional retardation in the ion migration. An explanation can be obtained if we invoke the relaxation effect, which is due to an emergence of an additional (retarding) field caused by a distortion of the EDL symmetry during the migration. Onsager showed that when $\varrho_{h}\ll 1$ the electrophoretic and relaxation terms are summed up [9], which in our notations yields
$1-\dfrac{K}{K_{0}}\simeq\varrho_{h}+\frac{1}{3}\left(1-\frac{1}{\sqrt{2}}\right)\frac{\ell_{B}}{\lambda_{D}}.$ (16)
Indeed, Eq. (16) provides an excellent fit to the data up to $R_{h}\sqrt{c_{\infty}}\simeq 6\times 10^{-3}$ (millimolar concentrations). Upon increasing concentration further the experimental data begin to approach to calculations from Eq. (14), and in concentrated solutions, $R_{h}\sqrt{c_{\infty}}\geq 8\times 10^{-2}$ (or $c_{\infty}\geq 0.5$ mol/l), our theory that includes an electrophoretic retardation solely is in excellent agreement with experiment. It is natural to conclude that when dilution is low the relaxation effect reduces with $c_{\infty}$ and disappears at high concentrations. Experimental observations [32, 33], simulations [6], and (numerical) calculations based on a nonlocal electrostatics approach [18] produce an endorsement of our conclusion. We also recall, that Overbeek [34] argued that, although in highly dilute solutions the relaxation effect grows with $c_{\infty}$ (see Eq. (16)), after reaching a maximum this correction to electrophoresis of colloid particles should reduce and disappears at high salt. Exactly the same was found here for inorganic salt solutions, which supports the concept of point-particle duality of ions.
IV Concluding remarks
[…]
  • The electrophoresis of inorganic ions can be accommodated within a theoretical framework that employs only a harmonic mean of hydrodynamic radii of ions and relies on a modified mean-field description of electrostatics (a concept of point-particle duality) and on a classical approach to colloid electrophoresis. The need to invoke specific constants as (arbitrary) parameters, or to correct an ionic concentration is thereby removed;
  • The conductivity of concentrated salt solutions is dictated solely by an ionic electrophoresis. Although the electrophoretic retardation dominates in highly dilute solutions too, to arrive at a more accurate description of their (in fact, extremely small) conductivity reduction, it seems necessary to include the relaxation correction.
This bears on the whole question of what we mean by a degree of dissociation that is often inferred from the conductivity measurements [36]. Our results show that the conductivity is interpreted without invoking the formation of ionic pairs, thus supporting the notion of complete dissociation of strong electrolytes. However remote from mainstream thinking this conclusion may seem, it would be worthwhile to recall that there are still lingering doubts about the reality of ion pairing, at least for univalent electrolytes in high permittivity solvents [5, 37].
Our considerations can be extended to asymmetric multivalent salts. The same concerns the temperature dependence of $K/K_{0}$, which follows from our theory, but requires the validation in terms of fit to experimental results. Another fruitful direction could be to consider the salt-dependence of a mobility of adsorbed ions [38, 39], which impacts electrokinetics [40].
On a more fundamental level, it would be interesting to revisit such issues as the diffusion of ions, effective permittivity and viscosity of salt solutions that can be quantitatively discussed using the standpoint taken here. If all these can be interpreted in a different from a common viewpoint way, the implications are large. A possible result would probably demand a revision of old dogmas and may represent a step forward.
Acknowledgements.
This research was supported by the Ministry of Science and Higher Education of the Russian Federation. We are indebted to E .S. Asmolov, R. Buchner, Y. A. Budkov, G. T. Hefter, B. Rotenberg, and G. A. Tsirlina for feedback and advises.

### v5

@@
 The conductivity of concentrated salt solutions has posed a real puzzle for theories of electrolytes.
-Despite a quantitative understanding of dilute solutions, an analytical theory for concentrated ones remains a challenge almost a century, although a number of parameters and effects incorporated into theories increases with time.
-Here we show that the conductivity of univalent salt solutions can be perfectly interpreted using a simple model that relies on a modified mean-field description of electrostatic effects and on a classical approach to calculating colloid electrophoresis.
-We derive a compact equation, which predicts that the ratio of conductivity at a finite concentration to that at an infinite dilution is the same for all salt, if it is plotted against a product of the harmonic mean of ion hydrodynamic radii and the square root of concentration.
+Despite a quantitative understanding of dilute solutions, an analytical theory for concentrated ones remains a challenge for almost a century, although a number of parameters and effects incorporated into theories increases with time.
+Here we show that the conductivity of univalent salt solutions can be perfectly interpreted using a simplest model that relies on a modified mean-field description of electrostatic interactions and on a classical approach to calculating colloid electrophoresis.
+We derive a compact equation, which predicts that the ratio of conductivity to that at an infinite dilution is the same for all salt and depends only on a product of the harmonic mean of ion hydrodynamic radii and the square root of concentration.
 Our equation fits very well the data for inorganic salts (up to a few mol/l), although at a very high dilution the relaxation correction seems necessary.
@@
 The physical origin of ionic conductivity is more or less clear.
-If an electric field $E$ is applied, ions migrate relative to a solvent by generating an electric current of density $J=KE$ (Ohm’s law).
-The migration speeds of ions are given by
+If a direct electric field $E$ is applied, ions migrate relative to a solvent with a constant speed towards one of the electrodes.
+Such a migration results in an electric current, which density obeys Ohm’s law, $J=KE$, provided the field is weak enough.
+The speeds of ions are given by
 $V_{\pm}=M_{\pm}E,$ (1)
@@
 The quantitative understanding of ion mobilities is a challenging problem that has been addressed over nearly a century and by many groups, which is often termed a central issue of chemical physics [6, 7].
-The simplest expression can be derived by postulating the Stokes resistance to the ion propulsion
+The simplest expression for univalent ions can be derived by postulating the Stokes resistance to the ion propulsion
 $M_{\pm}\simeq M_{\pm}^{0}\simeq\pm\dfrac{e}{6\pi\eta R_{\pm}},$ (3)
@@
 However, experiments on a conductivity show that $M$ and $K$ are generally smaller than predicted by (4) and this discrepancy augments on increasing salt concentration [8, 1].
-Chemists have long used this fact to infer a degree of dissociation (or an ion pairing).
-Physicists might view this simply as unreliability of hydrodynamic arguments based on the Stokes force.
+Chemists have long (since Arrhenius) used this fact to infer a degree of dissociation (or an ion pairing).
+Physicists might view this simply as unreliability of hydrodynamic arguments based on the Stokes drag force.
 There is a large literature describing attempts to provide a satisfactory theory of electrolyte conductivity.
-One of the first systematic treatments of the influence of $n_{\infty}$ on conductivity was contained in a remarkable paper by Onsager [9] who clarified a mechanism for the conductivity reduction, which involved ionic electrophoresis and relaxation, and showed that at a very high dilution the correction to $M_{0}$ depends on the square root of the concentration.
+One of the first systematic treatments of the influence of $n_{\infty}$ on conductivity was contained in a remarkable paper by Onsager [9] who argued that the relative conductivity reduction $K/K_{0}$ is attributed to two contributions.
+One (and the main) is ionic electrophoresis, which is the migration of ion surrounded by equilibrium screening cloud.
+Second is a so-called relaxation correction that depends on the cloud distortion produced by the motion in the external field.
+Onsager showed that at a very high dilution $K/K_{0}$ decreases linearly with the square root of the concentration.
 In efforts to better understand the connection between the ion mobilities and salt concentration many authors extended the theory, but failed to come to grips with the conductivity at high salt [10].
 However, most electrolyte solutions in nature and various applications are concentrated.
-Thus, salt concentration in human blood plasma is about 0.15 mol/l, in the Atlantic Ocean it is $\simeq 0.6$ mol/l, Li-ion batteries and supercapacitor electrolytes are usually of concentration $1-2$ mol/l, reference electrolytes of pH-meters and glass micropipette electrodes - of concentration 3 mol/l.
-Recent attempts at improvements on the theory has been focused on a more sophisticated description of electrostatic and hydrodynamic effects.
+Thus, salt concentration in human blood plasma is about 0.15 mol/l, in the Atlantic Ocean it is 0.6$\simeq$ 0.6≃ 0.6 mol/l, Li-ion batteries and supercapacitor electrolytes are usually of concentration 121-21 - 2 mol/l, reference electrolytes of pH-meters and glass micropipette electrodes - of concentration 3 mol/l.
+Recent attempts at improvements on the theory have been focused on a sophisticated description of so-called electrostatic and hydrodynamic effects based on microscopic concepts.
 Various techniques, such as the mean spherical approximation [11, 6], mode-coupling theory [12, 13], density functional approach [12, 14], and more [15, 16, 17, 7, 18] have been employed.
 These publications involve additional parameters and contributions, and mostly rely on numerical calculations.
-This makes them difficult to use and limits prediction capabilities, but simple analytical formulas applied for concentrated solutions have not yet been obtained, although there have been some attempts to deduce them from microscopic theories.
-Useful analytical solution is known [19], but expressed in terms of special functions and can hardly be called simple.
-Avni et al.
-[14] derived an approximate conductivity expression, but do not recognize its inapplicability at molar concentrations due to non physical behavior of electrophoretic and relaxation terms [20].
-Recently we suggested a completely different line to attack the ionic electrophoresis problem [21].
-Assuming cations and anions of equal size, we have carried calculations of an electrophoretic term and derived an equation for a corresponding reduction in relative conductivity, $K/K_{0}$, that applies even at a very low dilution.
-However, this paper has not given sufficient attention to the relaxation issue, which is traditionally invoked in the conductivity theories.
-It is also, in practice, unlikely that both cations and anions have exactly the same size, so the status of our theory for real salt solutions remains somewhat obscure.
-The present paper extends and generalizes the prior analysis [21] to solutions composed of inorganic ions of unequal radius $R_{+}\neq R_{-}$.
-We derive a simple equation for an electrophoretic contribution to conductivity valid up to a few mol/l and argue that relative conductivities, $K/K_{0}$, of all inorganic salts plotted against $R_{h}\sqrt{n_{\infty}}$ would collapse into a single curve.
+This makes them difficult to use and limits prediction capabilities, but simple analytical formulas that apply for concentrated solutions appear to be missing, although there have been some attempts to deduce them from microscopic theories.
+A useful analytical solution is known [19], but expressed in terms of special functions and can hardly be called simple or compact.
+A more recent approximate expression [14] is inapplicable at molar concentrations due to non physical behavior of both electrophoretic and relaxation terms [20].
+Thus, in spite of importance for many fields, a theory that has the merit of yielding useful (approximate) analytical result for the conductivity of concentrated salt solutions as well as makes strong appeal to macroscopic arguments is still absent.
+We also gained the impression that researchers in one branch appear to be unaware of closely related work in other branches of physics and chemistry.
+Indeed, most conductivity publications have either ignored or failed to make direct connection with the theory of electrophoresis developed in colloid science and fluid mechanics.
+Thus, already in 1921 Smoluchowski [21] made clear that the electrophoretic mobility is proportional to the electrostatic surface potential $\Phi_{s}$, and not, as one might imagine upon first thought, to the charge of the particle.
+Since in all existing theories ions are idealized as hard spheres of radius $a_{\pm}$, this conclusion holds for ions too, at least within the framework of such a colloid representation.
+However, the significance of the Smoluchowski paper does not seem to be recognized in the modern conductivity literature.
+Recently we have drawn attention to this drawback of the conductivity theories and have carried calculations of $\Phi_{s}$ and then of the so-called zeta potential of ions $Z\propto\Phi_{s}$, which defines their electrokinetic mobility [22].
+Assuming cations and anions of equal hydrodynamic radius, $R_{+}=R_{-}$, and using $a=R_{\pm}$, we have derived an equation for $K/K_{0}$, that applies even at a very low dilution and is more accurate and compact than prior approximate expression [14].
+However, our paper has not given attention to the relaxation issue, which is traditionally invoked in the conductivity theories.
+It is also, in practice, unlikely that both cations and anions have exactly the same size, so the status of the theory for real salt solutions still remains somewhat obscure.
+The present paper concerns with solutions composed of inorganic ions of unequal radius $R_{+}\neq R_{-}$.
+We revisit the issue of calculation of ion zeta potentials and propose an even much simpler expression than derived before [22], which appears to be more accurate and applies at a larger concentration range.
+We then use it to derive the simplest conductivity equation valid up to several mol/l, and argue that relative conductivities, $K/K_{0}$, of all inorganic salts plotted against $R_{h}\sqrt{n_{\infty}}$ would collapse into a single curve.
 This conclusion is supported by providing a comparison with data for several standard salts, which shows that our theory is in excellent agreement with experiment up to concentrations of a few molars.
@@
 II General considerations
-We consider a bulk 1:1 salt solution assuming that the description of its global static properties can be restricted to a mean-field Poisson-Boltzmann theory.
-The dimensionless electrostatic potential around each ion $\phi(z)=e\Phi(z)/(k_{B}T)$, where $k_{B}$ is the Boltzmann constant and $T$ is a temperature, represents a continuous function that depends on all other ions.
-The Debye screening length of a solution, $\lambda_{D}=\left(8\pi\ell_{B}n_{\infty}\right)^{-1/2}$, is defined as usually with the Bjerrum length, $\ell_{B}=\dfrac{e^{2}}{\varepsilon k_{B}T}$, where $\varepsilon$ is the solvent permittivity.
-By analyzing the experimental data it is more convenient to use the concentration $c_{\infty}[\rm{mol/l}]$, which is related to $n_{\infty}[\rm{m^{-3}}]$ as $n_{\infty}\simeq N_{A}\times 10^{3}\times c_{\infty}$, where $N_{A}$ is Avogadro’s number.
-The Bjerrum length of water at $T\simeq 298$ K is equal to about $0.7$ nm leading to a useful formula for 1:1 electrolyte
+Figure 1: Sketch of the charge distribution in the (a) real electrolyte solution, (b) Debye-Hückel classical model of finite size ions with the charges located at their center, and (c) model of ions with neutral interior, but constant surface charge densities.
+We consider a bulk 1:1 salt solution of concentration $n_{\infty}[\rm{m^{-3}}]$ assuming that the description of its global static properties can be restricted to a mean-field theory.
+Our approach is generally based on the fundamental ideas of Debye and Hückel [23].
+Namely, an electrolyte solution with randomly distributed ions [shown in Fig.
+1(a)] is idealized as a colloidal suspension of so-called central (positive and negative) ions representing nonconducting hard spheres of finite radii $a_{\pm}$ carrying unit charges ±$\pm e$ at their center.
+Compensating charges of the opposite sign and equal magnitude stay in the neighborhood of the central ions.
+These are the so-called ionic clouds or electrostatic diffuse layers (EDLs).
+EDLs are treated as composed by point-like ionic species, and their extension is of the order of the Debye screening length of a solution, $\lambda_{D}=\left(8\pi\ell_{B}n_{\infty}\right)^{-1/2}$, where $\ell_{B}=\dfrac{e^{2}}{\varepsilon k_{B}T}$ the Bjerrum length of the solvent, $\varepsilon$ is the solvent permittivity, $k_{B}$ is the Boltzmann constant, and $T$ is a temperature.
+This classical model is illustrated in Fig.
+1(b).
+Note that by analyzing the experimental data it is more convenient to use the concentration $c_{\infty}[\rm{mol/l}]$, which is related to $n_{\infty}$ as $n_{\infty}\simeq N_{A}\times 10^{3}\times c_{\infty}$, where $N_{A}$ is Avogadro’s number.
+The Bjerrum length of water at $T\simeq 298$ K is equal to about 0.70.70.7 nm leading to a useful formula for 1:1 electrolyte
 $\lambda_{D}[\rm{nm}]\simeq\frac{0.305[\rm{nm}]}{\sqrt{c_{\infty}[\rm{mol/l}]}}$ (5)
-Thus upon increasing $c_{\infty}$ from $10^{-5}$ to $3$ mol/l the screening length is reduced from about 100 down to 0.18 nm.
-Figure 1: Sketch of the cation and anion with hydrodynamic radii $R_{+}$ and $R_{-}$ in the bulk electrolyte solution characterized by $\lambda_{D}$.
-The cation propels with the speed $V_{+}$ in the direction of the electric field $E$.
-Its zeta potential $\zeta_{+}$ (or dimensionless mobility) is positive.
-The anion of negative $\zeta_{-}$ migrates against the field with the velocity $V_{-}$.
+Thus upon increasing $c_{\infty}$ from $10^{-5}$ to 444 mol/l the screening length (and the cloud extension) is reduced from about 100 down to 0.15 nm.
+Some further comments should be made.
+The fundamental insight of the Debye and Hückel approach is to realize that although the average potential inside the electrolyte solution is zero, there are strong positional correlations between the oppositely charged central ions [24].
+Indeed, if some cation is selected as central, then the position of a central anion cannot be arbitrary anymore, and their local ordering inevitably occurs, although on average they are randomly distributed.
+Another important insight is to realize what we would term the point-particle duality of ions (by analogy to the familiar concept of wave-particle duality of electrons in quantum mechanics): ions can be either point-like or exhibit properties of the particles of finite size depending on the circumstances.
+This classical model that is widely used to interpret equilibrium properties of electrolyte solutions will provide the starting point for our discussion of migration of central ions under an applied electric field.
+In the interior of the ions of radii $a_{\pm}$ there are no other charges except for the fixed at the centers, and the inner electrostatic potentials satisfy the Laplace equation, $\nabla^{2}\phi_{\pm}=0$, where $\phi_{\pm}=e\Phi_{\pm}/(k_{B}T)$ are the dimensionless electrostatic potentials.
+This inner field does not impact the electrophoresis, which depends only on the surface potentials $\phi_{s\pm}$ with respect to the solution far from them (with the zero average potential), as stressed in Sec.
+I.
+The distribution of electrostatic potentials around the central ions can be obtained by application of the non-linear Poisson-Boltzmann equation [25]:
+$\nabla^{2}\phi_{\pm}=\dfrac{1}{r^{2}}\dfrac{d}{dr}\left(r^{2}\dfrac{d\phi_{\pm}}{dr}\right)=\frac{\sinh\phi_{\pm}}{\lambda_{D}^{2}},$ (6)
+where $r$ is the distance from the centers of the corresponding spheres.
+The outer potentials would be the same as for the spheres with a constant surface charge density ±$\pm e/(4\pi a_{\pm}^{2})$.
+and decay from $\phi_{s\pm}$ to zero as $r\to\infty$, which is the first boundary condition that should be imposed to integrate (6).
+While this condition is applied at infinity, as we already discussed above, in reality the EDLs extend only to distances of the order of $\lambda_{D}$, and thus potentials in fact vanish beyond the clouds.
+The surface potentials are established self-consistently and salt-dependent.
+It is therefore admissible for electrophoretic calculations to replace the standard classical model shown in Fig.
+1(b) by spheres with a constant densities of a surface charge and neutral interior [see Fig.
+1(c)].
 Table 1: Some typical values of hydrodynamic radii of ions [26], they harmonic means and ratios.
@@
 LiI 0.238 0.1135 0.154 2.1
-We recall that the mean-field theory neglects correlations and finite sizes of ions.
-By contrast, to describe the dynamic response to an external field, the ion hydrodynamic radii $R_{\pm}$ should be taken into account.
+What is still missing and unknown are the radii $a_{\pm}$, which are the cut-off distances from the selected point charges, where there are no other ions.
+Some approaches to quantify them have been proposed [14, 6], but nevertheless, not developed enough to describe accurately the conductivity data.
+It is, however, well understood that to describe the dynamic response to an external field, the ion hydrodynamic radii $R_{\pm}$ should be taken into account.
 Inorganic ions have hydrodynamic radii from 0.1 to 0.3 nm [26], and we present their values for some univalent electrolytes in Table 1, together with harmonic means $R_{h}=2R_{+}R_{-}/(R_{+}+R_{-})$ and $N=R_{+}/R_{-}$.
-These apparent contradictions indicate that in reality, ions are neither point-like nor particles of finite size, exhibiting the properties of the former or the latter only depending on the circumstances.
-This can be referred to as the point-particle duality of ions (by analogy to the familiar concept of wave-particle duality of electrons in quantum mechanics).
-Such a duality can be introduced, if an electrolyte solution is modelled as a colloidal suspension of so-called central ions representing isolating particles of radius $R_{\pm}$ with a constant surface charge density ±$\pm e/(4\pi R_{\pm}^{2})$ [21].
-Such a sphere induces the same outer field as a point charge ±$\pm e$ located at its center would, and the salt-dependent surface potential $\phi_{s\pm}$ is established self-consistently.
-By contrast, the electrostatic diffuse layers (EDLs or clouds) are treated on the classical mean-field level and formed by point-like ionic species, thus their extension is defined by $\lambda_{D}$.
-Two further comments should be made: (i) Since the model prevents point-like EDL ions from penetrating inside the sphere, the need to invoke an additional electrostatic cut-off [14] is removed.
-(ii) The approach neglects electrostatic correlations inside the cloud, but provides positional correlations between central ions, although on average they are randomly distributed.
+Since the introduction of $R_{\pm}$ inevitably prevents point-like EDL ions from penetrating inside the sphere, the need to invoke an additional electrostatic cut-off [14] is removed.
+So, below we simply postulate $a_{\pm}=R_{\pm}$ and impose the electrostatic boundary condition $\phi^{\prime}_{\pm}(r=R_{\pm})=\mp\ell_{B}/R_{\pm}^{2}$ to integrate Eq.
+(6) at this surface.
 Since the electrostatics of ions involves two length scales, it is convenient to introduce their dimensionless radii $\varrho_{\pm}=R_{\pm}/\lambda_{D}$.
-Note that in the range of $c_{\infty}$ below 3 mol/l we consider here, the values of ϱ±subscriptitalic-$\varrho_{\pm}$ remain smaller than unity or very close to it.
-Say, if we set $c_{\infty}=2$ mol/l, then for a largest ion (Li+) in Table 1 we obtain $\varrho_{+}\simeq 1$, but for a smallest (I-) we get $\varrho_{-}\simeq 0.5$.
-When an electric field $E$ is applied, an electro-osmotic flow around central ions, referred below simply to as ions, is induced.
-The electroosmosis takes its origin in the EDL, where a tangential electric field generates a force that sets the fluid in motion.
-The emergence of this flow in turn provides hydrodynamic stresses that cause the propulsion of ions with a velocity given by Eq.(1) as sketched in Fig.
-1.
-The space between clouds of two ions is electro-neutral, and this region is force-free (no flow) provided they do not overlap.
-Thus, the problem of ionic electrophoresis can be treated on a single-particle level.
-We return to the importance of the relaxation later, by focusing now on electrophoresis solely.
-This contribution to a conductivity is traditionally calculated assuming that the surrounding cloud remains undisturbed during the migration.
-The electrophoretic mobilities of these ions can be defined using the known expression for an electro-osmotic mobility [23, 24] taken with the opposite sign
+Note that in the range of $c_{\infty}$ below 2 mol/l the values of $\varrho_{\pm}$ remain smaller than unity or very close to it.
+Say, if we set $c_{\infty}=2$ mol/l, then for the largest ion (Li+) in Table 1 we obtain $\varrho_{+}\simeq 1.1$, but for the smallest (I-) we get $\varrho_{-}\simeq 0.5$.
+However, when $c_{\infty}=4$ mol/l the same ions give $\varrho_{+}\simeq 1.6$ and $\varrho_{-}\simeq 0.7$.
+Figure 2: Sketch of the cation and anion with hydrodynamic radii $R_{+}$ and $R_{-}$ in the bulk electrolyte solution characterized by $\lambda_{D}$.
+The cation propels with the speed $V_{+}$ in the direction of the electric field $E$.
+Its zeta potential $\zeta_{+}$ (or dimensionless mobility) is positive.
+The anion of negative $\zeta_{-}$ migrates against the field with the velocity $V_{-}$.
+When an electric field $E$ is applied, an electro-osmotic flow around central anions and cations, referred below simply to as ions, is induced.
+The electroosmosis takes its origin in their EDLs, where a tangential electric field generates a force that sets the fluid in motion.
+The emergence of this flow in turn provides hydrodynamic stresses that cause the propulsion of ions with a velocity given by Eq.
+(1) as sketched in Fig.
+2.
+If the clouds of two ions do not overlap, the space between EDLs remains electro-neutral, and, therefore, this region is force-free.
+Consequently, in this outer region there is no fluid flow.
+Thus, within our model the problem of ionic electrophoresis can be treated on a single-particle level provided the concentration is not too high (below 4 mol/l or so).
+This, of course, significantly simplifies the problem.
+However, it would be worthwhile to stress that even the problem for a single colloidal sphere, the so-called “standard electrokinetic model” [27], can in general be solved only numerically [since analytical solution of Eq.
+(6) has not been found yet].
+In the standard considerations of the electrophoretic migration, an essential supposition is normally that the radial electric field due to the double layer and the externally applied field may be simply superimposed on each other [28].
+The Péclet number is assumed to be small, so that the convective fluxes of ions are ignored.
+If so, the system of electrokinetic equations is decoupled and the potential obeys the Poisson-Boltzmann equation [29], implying that the EDLs are undisturbed by the migration.
+Such a situation will be referred below simply to as electrophoresis.
+This is certainly not always the case, as the EDL symmetry may be distorted during the migration.
+This effect that is termed relaxation gives rise to an additional (retarding) field, which is directed oppositely to $E$.
+Note that (as a second order effect) the relaxation retardation, if any, is smaller than the one, caused by the motion of the ions of the undisturbed cloud.
+Since Onsager it is accepted that to arrive at a complete description of migration of ions, it is necessary to include this relaxation effect [9].
+Note, however, the correction for relaxation goes to zero for large colloidal particles and is never included into consideration [28].
+III Results and discussion
+We return to the issue of importance of the relaxation later, by focusing first on electrophoresis solely.
+Recall that this (leading order) contribution to a conductivity is traditionally calculated assuming that the clouds remain undisturbed during the migration.
+The electrophoretic mobilities of ions can be defined using the known expression for an electro-osmotic mobility [21, 30] taken with the opposite sign
 $M_{\pm}=\dfrac{\varepsilon Z_{\pm}}{4\pi\eta}=\dfrac{e}{4\pi\eta\ell_{B}}\zeta_{\pm},$ (7)
@@
 In such a definition the geometry factor is hidden inside $\zeta$.
-Note that by introducing the dimensional velocity $v_{\pm}=\dfrac{4\pi\eta\ell_{B}}{eE}V_{\pm}$, one can rewrite (1) as $v_{\pm}=\zeta_{\pm},$ which points clearly that the electrophoretic speed of ions is set by their zeta potentials.
-III Results and discussion
-Zeta potentials of small inorganic ions are given by [21]
+Note that by introducing the dimensionless velocity $v_{\pm}=\dfrac{4\pi\eta\ell_{B}}{eE}V_{\pm}$, one can rewrite (1) as $v_{\pm}=\zeta_{\pm},$ which points clearly that the electrophoretic speed of ions is set by their zeta potentials.
+The expression for zeta potentials of spherical isolating particles, which obey hydrodynamic no-slip boundary condition, has been derived already by Henry [31], and for ions of radius $R_{\pm}$ this can be formulated as
+[…]
+In order to take the integrals in Eq.
+(8) it is necessary to introduce explicitly the dependence of potentials on $r$.
+The $\phi_{\pm}$-profiles (and, consequently, $\phi_{s\pm}$) can be computed from Eq.
+(6) leading then to a numerical solution for $\zeta_{\pm}$ of ions.
+When the linear approximation of Eq. (6) can be applied the local and surface potentials are given, approximately, by
+$\phi_{\pm}\simeq\pm\dfrac{\ell_{B}}{\left(1+\varrho_{\pm}\right)}\dfrac{\mathrm{e}^{\varrho_{\pm}-r/\lambda_{D}}}{r},\,\phi_{s\pm}\simeq\pm\dfrac{\ell_{B}}{R_{\pm}\left(1+\varrho_{\pm}\right)}.$ (9)
+It is normally considered that this is justified provided $|\phi_{s\pm}|\leq 1$, but we have recently shown that if radii are small, the linearization of Eq.
+(6) becomes no longer specific just to low potentials [22].
+Thus Eqs.
+(9) should be a sensible approximation for inorganic salt ions of $|\phi_{s\pm}|=O(1)$.
+Substituting equation (9) into (8) and performing the integration yields
 $\zeta_{\pm}\simeq\phi_{s\pm}\mathcal{F_{\pm}}\simeq\pm\dfrac{\ell_{B}}{R_{\pm}(1+\varrho_{\pm})}\mathcal{F_{\pm}},$ (10)
-where $\mathcal{F_{\pm}}$ are the special functions derived by Henry [25]
+where $\mathcal{F_{\pm}}$ are the Henry special functions for nonconducting spherical particles [31]
 ±$\mathcal{F_{\pm}}=1-e^{\varrho_{\pm}}\left[5\operatorname{Ei}_{7}(\varrho_{\pm})-2\operatorname{Ei}_{5}(\varrho_{\pm})\right].$ (11)
 […]
-±$\mathcal{F_{\pm}}\simeq\dfrac{2}{3}\left[1+\left(\dfrac{\varrho_{\pm}}{4}\right)^{2}\right]$ (9)
-[…]
-$\zeta_{\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}(1+\varrho_{\pm})}.$ (10)
-Clearly, the upper bound for the zeta potential is attained when $\rho_{\pm}\to 0$ (an infinite dilution)
+Figure 3: Function ±$\mathcal{F_{\pm}}/(1+\rho_{\pm})$ vs $\rho_{\pm}$ with $\mathcal{F_{\pm}}$ calculated from Eqs.
+(11) (solid curve), (12) (dashed curve), and ±$\mathcal{F_{\pm}}=2/3$ (dash-dotted curve).
+Dotted curve shows the electrophoretic term derived in [14].
+From Eqs.
+(10) and (11) if follows that the salt dependence of $\zeta_{\pm}$ is hidden only in ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$.
+This dependence is illustrated in Fig.
+3, where we plot ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ as a function of $\varrho_{\pm}$.
+The calculations are first made using $\mathcal{F_{\pm}}$ given by Eq.
+(11).
+It can be seen that the function ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ in Eq.
+(10) decreases [and vanishes in the limit $\lambda_{D}\to 0$].
+Thus on increasing $\varrho_{\pm}$ (reducing $\lambda_{D}$) the migration speed of ions reduces, which is self-evident since the electrophoresis takes its origin in the EDL.
+For small $\varrho_{\pm}$ Eq. (11) can be expanded about $\varrho_{\pm}=0$ and reduces to
+±$\mathcal{F_{\pm}}\simeq\dfrac{2}{3}\left[1+\left(\dfrac{\varrho_{\pm}}{4}\right)^{2}\right],$ (12)
+which has been used to derive the approximate conductivity equation [22].
+To examine its range of applicability the curve for ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ calculated with $\mathcal{F_{\pm}}$ given by Eq.
+(12) is included in Fig.
+3.
+The accuracy is quite good for $\varrho_{\pm}\leq 1$, but at larger salt concentrations there is some discrepancy in the direction of larger ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ than predicted by using Eq.
+(11).
+Moreover, the function exhibits its minimum at $\varrho_{\pm}\simeq 3$ and then augments, although can only decrease.
+Note, however, that the electrophoretic term derived by Avni et al.
+[14],
+±$\frac{\mathcal{F_{\pm}}}{1+\varrho_{\pm}}\simeq\frac{2}{3}\left(1-\varrho_{\pm}e^{-a/\lambda_{D}}\right),$ (13)
+where $a$ is the is the cut-off length defined as the sum of the cation and anion crystallographic radii, fails at much smaller $\varrho_{\pm}$.
+The calculation from Eq.
+(13) is included in Fig.
+3.
+It can be seen that it shows significant deviations [from ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ obtained using (11)] already at $\varrho_{\pm}\geq 2\times 10^{-1}$ and turns to 2/32/32 / 3 when $\varrho_{\pm}\geq 2$.
+This implies that the zeta potentials become the same as at an infinite dilution, which is counterintuitive.
+Also included in Fig.
+3 is the curve calculated using ±$\mathcal{F_{\pm}}=2/3$.
+It seems clear that this provides ±$\mathcal{F_{\pm}}/(1+\varrho_{\pm})$ that practically coincides (when $\varrho_{\pm}\leq 1$) or is very close [at $\varrho_{\pm}=O(1)$] to obtained with Eq.
+(11).
+This implies that the decrease in $|\zeta_{\pm}|$ with salt is to the first order caused only by the reduction of the ion surface potential:
+$\zeta_{\pm}\simeq\frac{2}{3}\phi_{s\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}(1+\varrho_{\pm})}.$ (14)
+The upper bound for the zeta potential is attained when $\rho_{\pm}\to 0$ (an infinite dilution)
 $\zeta_{\pm}\to\zeta_{0\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}}.$ (15)
-Figure 2: Zeta potentials $\zeta_{\pm}$ computed for K+, Li+, Cl-, I- (solid curves from top to bottom). Dashed lines are obtained using Eq. (11). Symbols show calculations from Eq. (10).
-Figure 2 includes computed zeta potentials of several ions.
-For these examples we take from Table 1 cations (K+, Li+) and anions (I-, Cl-) of smallest and largest radius.
-Calculations are made using the first equality in (7) with $\phi_{s\pm}$ found numerically from the nonlinear Poisson-Boltzmann equation and $\mathcal{F_{\pm}}$ calculated from Eq.
+Figure 4: Zeta potentials $\zeta_{\pm}$ vs.
+$c_{\infty}$ computed for K+, Li+, Cl-, I- (solid curves from top to bottom).
+Dashed lines are obtained using Eq.
 (15).
-The straight lines corresponding to $\zeta_{0\pm}$ given by (11) describe perfectly the distinct plateau regions at very low concentrations.
+Symbols show calculations from Eq.
+(14).
+Figure 4 includes computed zeta potentials of several ions as a function of salt concentration.
+For these examples we take from Table 1 cations (K+, Li+) and anions (I-, Cl-) of the smallest and largest radius.
+As a side note, these numerical calculations [based on Eqs.
+(8) and (6)] fully coincide with those made from the first equality in (10) using $\phi_{s\pm}$ computed from the nonlinear Poisson-Boltzmann equation and $\mathcal{F_{\pm}}$ given by Eq.
+(11).
+The straight lines corresponding to $\zeta_{0\pm}$ given by (15) describe perfectly the distinct plateau regions at low concentrations.
 On increasing concentration, however, the absolute values of $\zeta_{\pm}$ decrease.
-Also included are calculations from (10).
+Also included are calculations from Eq.
+(14).
 It can be seen that their fit to numerical data is extremely good in the whole concentration range.
-Refer to captionRefer to captionRefer to caption
-Figure 3: $K_{\infty}/K_{\infty}^{0}$ as a function of $c_{\infty}$ for KBr, NaCl and LiI aqueous solution calculated from Eqs.
+Figure 5: $K/K_{0}$ as a function of $c_{\infty}$ for KBr, NaCl and LiI aqueous solution calculated from Eqs.
 (17), (18) and (19) (solid, dash-dotted and dashed curves).
@@
 $\dfrac{M}{M_{0}}\equiv\dfrac{K}{K_{0}}\simeq\dfrac{1}{N+1}\left(\dfrac{1}{1+\varrho_{+}}+\dfrac{N^{2}}{N+\varrho_{+}}\right),$ (17)
-where $M_{0}$ and $K_{0}$ are given by Eqs.
-(4).
+where $M_{0}$ and $K_{0}$ are given by Eqs. (4).
 Figure 5 shows $K/K_{0}$ calculated from Eq.
 (17) as a function of $c_{\infty}$.
-The calculations are compared with experimental data for three standard salts [27].
+The calculations are compared with experimental data for three standard salts [33, 34, 35, 36].
 As seen from Table 1, for these examples $N$ varies from 1.1 for KBr, but can be larger than 2 (e.g.
@@
 An overall conclusion from this linear scale plot is that the theoretical curve for $K/K_{0}$ is well consistent with experiment.
-At large concentrations Eq.(13) predicts either slightly lower (KBr), but still fits data quite well.
-When $N=1$, we get $\varrho_{h}=R_{h}/\lambda_{D}=2\varrho_{+}/(1+N)=\varrho_{+}$, and Eq.
-(13) reduces to
+Only for KBr Eq.(17) predicts a slightly lower conductivity at large concentrations, but still fits data quite well.
+This is a startling result since our mean-field theory omits nonidealities that are normally considered to be important at high salt [and at this point does not include the relaxation effect].
+Thus, our results suggest that, if any, they constitute a second-order correction, which could make only a small improvement to the data description.
+When $N=1$, we get $\varrho_{h}=R_{h}/\lambda_{D}=2\varrho_{+}/(1+N)=\varrho_{+}$, and Eq. (17) reduces to
 $\dfrac{K}{K_{0}}\simeq\dfrac{1}{1+\varrho_{h}}.$ (18)
 Eq.
-(14) predicts a monotonic decrease in relative (electrophoretic) conductivity with $\varrho$ as it should be, and $K/K_{0}\to 0$ as $\varrho\to\infty$.
+(18) predicts a monotonic decrease in relative (electrophoretic) conductivity with $\varrho_{h}$ as it should be, and $K/K_{0}\to 0$ as $\varrho_{h}\to\infty$.
 For $\varrho_{h}\ll 1$ it reduces to the famous Onsager formula for the electrophoretic effect
 $\dfrac{K}{K_{0}}\simeq 1-\varrho_{h}.$ (19)
-The calculation from Eq.(14) is also included in Fig.
-3.
+The calculation from Eq.
+(18) is also included in Fig.
+5.
 It can be seen that deviations from Eq.
 (17) are extremely small.
-The difference between (13) and (14) is always below a few %.
+The difference between (17) and (18) is always below a few percent.
 We thus conclude that compact approximate Eq.
@@
 That this is indeed so is demonstrated in Fig.
-4, where we plot in a semi-log scale a relative conductivity calculated from (14) vs.
-$R_{h}\sqrt{c_{\infty}}$ along with the experimental data for a variety of inorganic salts at $T=298$[27, 29, 30, 31, 28].
+6, where we plot in a semi-log scale a relative conductivity calculated from (18) vs.
+$R_{h}\sqrt{c_{\infty}}$ along with the experimental data for a variety of inorganic salts at $T=298\leavevmode\nobreak$[33, 35, 36, 37, 34].
 Besides salts presented in Table 1 we include KCl, LiCl, and LiClO4 using their $R_{h}$ = 0.127, 0.163, and 0.174 nm, correspondingly, to re-scale the concentration [26].
-Note that since the theory is justified provided $R_{h}\sqrt{c_{\infty}}$ is below ca.
-0.305, we employ for this plot only data for $c_{\infty}\leq 3$ mol/l.
-One can see that the approximate theory is in a good agreement with the data for all salts, perhaps with only insignificant discrepancy at low dilution that is probably comparable with the experimental error.
-The discrepancy is always in the direction of smaller $K/K_{0}$ than predicted by Eq.
-(14), but the data for KBr and KCl show slightly larger $K/K_{0}$.
+Note that we employ for this plot data for $c_{\infty}$ up to 4 mol/l.
+One can see that the approximate theory is in a good agreement with the data for all salts, perhaps with only insignificant discrepancy at intermediate dilution.
+The discrepancy is always in the direction of slightly smaller $K/K_{0}$ than predicted by Eq.
+(18).
+However, at low dilution our theoretical calculations practically coincide with the experimental results.
+Only the data for KBr and KCl show slightly larger $K/K_{0}$.
+Note that there exist some data [38] suggesting that conductivity of potassium salts is even higher than in specimen examples used here, but not supported by other experiments.
 The reason for qualitative differences only for these two salts is unclear and is open for discussion, but it must be remembered that this was a first-order calculation only, and we did not expect it to be very accurate.
@@
 (19) is also included in Fig.
-4.
+6.
 Clearly, the data obtained at high salt are irreconcilable with this linear equation.
@@
 Symbols show the same data as in Fig.
-4.
+6.
 Dashed-dotted, dashed and dotted lines are calculated from Eqs.
@@
 6 are reproduced in Fig.
-5, where a decrement to a relative conductivity, $1-K/K_{0}$, is plotted against $R_{h}\sqrt{c_{\infty}}$ in a log-log scale.
+7, where a decrement to a relative conductivity, $1-K/K_{0}$, is plotted against $R_{h}\sqrt{c_{\infty}}$ in a log-log scale.
 It can be seen that in dilute solutions, where the decrease in relative conductivity is so small that was even not discernible on the scale of Fig.
-4, the data are slightly above the theory indicating some additional retardation in the ion migration.
-An explanation can be obtained if we invoke the relaxation effect, which is due to an emergence of an additional (retarding) field caused by a distortion of the EDL symmetry during the migration.
-Onsager showed that when $\varrho_{h}\ll 1$ the electrophoretic and relaxation terms are summed up [9], which in our notations yields
+6, the data are slightly above the theory indicating some additional retardation in the ion migration.
+The discrepancy looks constant at smaller values $R_{h}\sqrt{c_{\infty}}$, which indicates an increase in the extra retardation effect with salt, but then the deviations from our (electrophoretic) theory reduce and completely disappear in concentrated solutions.
+An explanation can be obtained if we invoke the relaxation effect.
+Onsager showed that when $\varrho_{h}\ll 1$ the electrophoretic and relaxation terms can be considered separately and are summed up [9], which in our notations yields
 $1-\dfrac{K}{K_{0}}\simeq\varrho_{h}+\frac{1}{3}\left(1-\frac{1}{\sqrt{2}}\right)\frac{\ell_{B}}{\lambda_{D}}.$ (20)
-Indeed, Eq.
-(16) provides an excellent fit to the data up to $R_{h}\sqrt{c_{\infty}}\simeq 6\times 10^{-3}$ (millimolar concentrations).
-Upon increasing concentration further the experimental data begin to approach to calculations from Eq.
-(14), and in concentrated solutions, $R_{h}\sqrt{c_{\infty}}\geq 8\times 10^{-2}$ (or $c_{\infty}\geq 0.5$ mol/l), our theory that includes an electrophoretic retardation solely is in excellent agreement with experiment.
-It is natural to conclude that when dilution is low the relaxation effect reduces with $c_{\infty}$ and disappears at high concentrations.
-Experimental observations [32, 33], simulations [6], and (numerical) calculations based on a nonlocal electrostatics approach [18] produce an endorsement of our conclusion.
-We also recall, that Overbeek [34] argued that, although in highly dilute solutions the relaxation effect grows with $c_{\infty}$ (see Eq.
-(16)), after reaching a maximum this correction to electrophoresis of colloid particles should reduce and disappears at high salt.
-Exactly the same was found here for inorganic salt solutions, which supports the concept of point-particle duality of ions.
+Indeed, Eq. (20) provides an excellent fit to the data up to $R_{h}\sqrt{c_{\infty}}\simeq 6\times 10^{-3}$ (millimolar concentrations) as displayed in Fig. 7.
+Upon increasing concentration further the experimental data begin to approach the results of calculations from Eq.
+(18), and in concentrated solutions, $R_{h}\sqrt{c_{\infty}}\geq 8\times 10^{-2}$ (or $c_{\infty}\geq 0.5$ mol/l), the theory that includes an electrophoresis solely is in excellent agreement with experiment.
+It is natural to conclude that, although in highly dilute solutions the relaxation effect grows with $c_{\infty}$ [see Eq.
+(20)], after reaching a maximum this correction to electrophoresis of inorganic salt ions should reduce and disappears at high salt.
+As the side note, exactly the same trends have been predicted already by Overbeek [28] for standard colloid particles.
+Experimental observations [39, 40], integral equations [6], and (numerical) calculations based on a nonlocal electrostatics approach [18] support our conclusion about disappearance of relaxation at high salt.
+In any event, it seems even intuitively clear that the relaxation, which is caused by an emergence of an additional retarding field due to a distortion of the cloud symmetry, should vanish at very high salt: no EDLs - nothing to distort, and (as a second order effect) is always smaller than electrophoretic effect.
+However, the magnitude of the relaxation term in the approximate conductivity equation derived by Avni et al.
+[14] increases monotonically with salt, as highlighted in recent publications [6, 20].
+When $a=O(\lambda_{D})$ this term becomes of the same order as electrophoretic, and in more concentrated solutions even dominates.
+As a result, when the EDL nearly disappears its relaxation grows.
+Clearly, it cannot be so, and in essence, the equation by Avni et al.
+[14] does not apply at molar concentrations [41].
+However, simple conductivity equation (18), which provides a physically correct asymptotic behavior of the electrophoretic term and neglects relaxation, is very well suited to describe concentrated salt solutions.
 IV Concluding remarks
 […]
-  • The electrophoresis of inorganic ions can be accommodated within a theoretical framework that employs only a harmonic mean of hydrodynamic radii of ions and relies on a modified mean-field description of electrostatics (a concept of point-particle duality) and on a classical approach to colloid electrophoresis.
+  • The electrophoresis of inorganic ions can be accommodated within a theoretical framework that employs only a harmonic mean of hydrodynamic radii of ions and relies on a mean-field description of electrostatics and on a classical approach to colloid electrophoresis.
 The need to invoke specific constants as (arbitrary) parameters, or to correct an ionic concentration is thereby removed;
-  • The conductivity of concentrated salt solutions is dictated solely by an ionic electrophoresis.
+  • The conductivity of concentrated salt solutions is dictated solely by ionic electrophoresis.
 Although the electrophoretic retardation dominates in highly dilute solutions too, to arrive at a more accurate description of their (in fact, extremely small) conductivity reduction, it seems necessary to include the relaxation correction.
-This bears on the whole question of what we mean by a degree of dissociation that is often inferred from the conductivity measurements [36].
+This bears on the whole question of what we mean by a degree of dissociation that is often inferred by chemists from the conductivity measurements [42].
 Our results show that the conductivity is interpreted without invoking the formation of ionic pairs, thus supporting the notion of complete dissociation of strong electrolytes.
-However remote from mainstream thinking this conclusion may seem, it would be worthwhile to recall that there are still lingering doubts about the reality of ion pairing, at least for univalent electrolytes in high permittivity solvents [5, 37].
+However remote from mainstream thinking this conclusion may seem, it would be worthwhile to recall that there are still lingering doubts about the reality of ion pairing, at least for univalent electrolytes in high permittivity solvents [5, 24, 43].
 Our considerations can be extended to asymmetric multivalent salts.
@@
 This research was supported by the Ministry of Science and Higher Education of the Russian Federation.
-We are indebted to E .S.
+We are indebted to E.
+S.
 Asmolov, R.
-Buchner, Y.
-A.
-Budkov, G.
+Buchner, G.
 T.
@@
 A.
-Tsirlina for feedback and advises.
+Tsirlina for feedback and advice.
+DATA AVAILABILITY
+The data that support the findings of this study are available within the article.
+AUTHOR DECLARATIONS
+The authors have no conflicts to disclose.